OpenStudy (anonymous):

Assignment: Add and Subtract Rational Expressions Choose any four (4) of the following five problems to solve. Show all work leading to your answer. 1. Randy drives his car to work each day. Suppose the expression 6 6 x x3 represents the total time it takes Randy to drive to work in a day, where x represents the average speed of the car in miles per hour, during the first part of the drive. a. Use addition to simplify the expression 6 6 x x3 . Show all the steps necessary to write the expression with a common denominator and to simplify the expression.

OpenStudy (ja1):

wait what is 6 6 x x3 ???

OpenStudy (anonymous):

6/x+6/3x sorry

OpenStudy (ja1):

Oh ok... well I don't know how to solve that xD

OpenStudy (anonymous):

oh ok .

OpenStudy (ja1):

sowwy :/

OpenStudy (anonymous):

it's ok :) thanks for trying to help me

OpenStudy (irishboy123):

the second fraction is 6/3x. you can start by simplifying that. its the same as 2/x, isn't it? just divide the top AND bottom of the fraction by 3, meaning the ratio of the numbers stays the same. think 1/2 = 2/4 = 4/8 = 8/16. then your sum is simple, you are adding 6/x to 2/x. you are adding 1/x ths together. you can add 1/3 rds to 1/4 ths, you need a common denominator. here the common denominator is x! you should get 8/x. is that clear? please advise if not.